# Rigged Dynamic Mode Decomposition: Data-Driven Generalized Eigenfunction Decompositions for Koopman Operators

### v1

@@ Abstract
We introduce the Rigged Dynamic Mode Decomposition (Rigged DMD) algorithm, which computes generalized eigenfunction decompositions of Koopman operators. By considering the evolution of observables, Koopman operators transform complex nonlinear dynamics into a linear framework suitable for spectral analysis. While powerful, traditional Dynamic Mode Decomposition (DMD) techniques often struggle with continuous spectra. Rigged DMD addresses these challenges with a data-driven methodology that approximates the Koopman operator’s resolvent and its generalized eigenfunctions using snapshot data from the system’s evolution. At its core, Rigged DMD builds wave-packet approximations for generalized Koopman eigenfunctions and modes by integrating Measure-Preserving Extended Dynamic Mode Decomposition with high-order kernels for smoothing. This provides a robust decomposition encompassing both discrete and continuous spectral elements. We derive explicit high-order convergence theorems for generalized eigenfunctions and spectral measures. Additionally, we propose a novel framework for constructing rigged Hilbert spaces using time-delay embedding, significantly extending the algorithm’s applicability. We provide examples, including systems with a Lebesgue spectrum, integrable Hamiltonian systems, the Lorenz system, and a high-Reynolds number lid-driven flow in a two-dimensional square cavity, demonstrating Rigged DMD’s convergence, efficiency, and versatility. This work paves the way for future research and applications of decompositions with continuous spectra.
@@ 1 Introduction
This paper introduces a new Koopman mode decomposition and computational framework for unitary Koopman operators on a rigged Hilbert space (see Section 2.4). In the rigged Hilbert space, the discrete and continuous spectrum are treated on equal footing, and (almost) every point in the spectrum is associated with a set of coherent generalized eigenfunctions. Generalized eigenfunctions associated with the continuous spectrum are not elements of the Hilbert space. Instead, as distributions in the rigged Hilbert space, they represent coherent linear features of the observables that may be used to identify coherent patterns in Eq. 1.
Our algorithm, which we call Rigged DMD, computes smooth wave-packet approximations to generalized Koopman eigenfunctions using a two-stage framework (see Fig. 1). In the first stage, we use mpEDMD to construct a finite-dimensional unitary approximation to the Koopman operator from the snapshot data in Eq. 2. In the second stage, we use the mpEDMD resolvent to construct wave packets of observables with highly localized spectral content. These approximations converge to generalized eigenfunctions associated with both the discrete and the continuous spectrum (see Section 5). They are approximately coherent (see Theorem 5.13), achieve high orders of accuracy under natural regularity conditions on the spectral measure of $\mathcal{K}$ (see Theorem 5.9), and converge pointwise in regions where generalized eigenfunctions can be identified with smooth functions (see Theorem 5.11). For example, Figure 2 shows four wave-packet approximations to generalized eigenfunctions associated with the continuous spectrum of a Koopman operator on the Lorenz attractor.
There is extensive literature on rigged Hilbert spaces for self-adjoint operators in PDE analysis and quantum mechanics [40, 14, 49, 13, 12]. However, standard rigging techniques do not always apply to Koopman operators (consider the task of constructing a nuclear space on the Lorenz attractor). For this reason, we provide a novel method to construct nuclear spaces for unitary operators with sparse, infinite-dimensional matrix representations (see Section 6). This technique can be applied to any Koopman operator via time-delay embedding, enabling the computation of generalized Koopman eigenfunctions for the Lorenz attractor.
@@ 2 Background: Koopman mode decompositions
Such systems preserve an energy or volume associated with $\omega$ and are ubiquitous [7, 34, 74, 46, 86]. Many systems admit invariant measures [55] or measure-preserving post-transient behavior [61]. We also assume that the dynamical system is invertible modulo $\omega$-null sets (see [35, Chapter 7]), so that $\mathcal{K}$ is a unitary operator on the Hilbert space $L^{2}(\Omega,\omega)$:
@@ 2.2 Spectrum and pseudoeigenfunctions
Since $\mathcal{K}$ is unitary, we have $\mathrm{Sp}(\mathcal{K})\subset\mathbb{T}=\{z\in\mathbb{C}:|z|=1\}$. The spectrum includes the set of eigenvalues of $\mathcal{K}$, but in general, it may contain points that are not eigenvalues and hence do not have an associated eigenfunction. However, there are approximately coherent observables associated with each point because $\mathcal{K}$ is unitary (therefore, normal). For each point $\lambda\in\mathrm{Sp}(\mathcal{K})$ and any $\delta>0$, there exists a $\phi_{\delta}\in L^{2}(\Omega,\omega)$ with norm one such that $\|(\mathcal{K}-\lambda I)\phi_{\delta}\|\leq\delta$. The observable $\phi_{\delta}$ is called a $\delta$-pseudoeigenfunction [81, p. 31] and satisfies
@@ 2.3 Spectral measures and diagonalization
The spectral theorem [29, Thm. X.4.11] provides a diagonalization of $\mathcal{K}$ using a projection-valued measure $\mathcal{E}$ supported on $\mathrm{Sp}(\mathcal{K})\subset\mathbb{T}$. To each Borel subset $S\subset\mathbb{T}$, $\mathcal{E}$ associates an orthogonal projection $\mathcal{E}(S)$, which one should think of as the spectral content of $\mathcal{K}$ on $S$. In particular, we have the following decompositions:
@@ 2.4 Generalized eigenfunctions
To achieve a modal decomposition using generalized eigenfunctions, we require that $\mathcal{S}$ also be a countably Hilbert nuclear space [41, Ch. 1.3]. Then, there is a complete set of generalized eigenfunctions $\{g_{\theta,k}:\theta\in{[{-}\pi,\pi]_{\mathrm{per}}},k\in\mathcal{I}\}\subset\mathcal{S}^{*}$ such that for any $g\in\mathcal{S}$, it holds in $L^{2}(\Omega,\omega)$:
(10) $g=\sum_{k\in\mathcal{I}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!\langle g_{\varphi,k}^{*}|g\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi),\quad\mathcal{K}g=\sum_{k\in\mathcal{I}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!\exp(i\varphi)\langle g_{\varphi,k}^{*}|g\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi).$
The spectral measures $\{\xi_{g_{k}}\}$ reflect a choice of normalization and the multiplicity of $\lambda=\exp(i\theta)\in{\rm Sp}(\mathcal{K})$ is the number of nonzero functionals $\{g_{\theta,k}\}$. This expansion replaces the projection-valued measure expansion in Eq. 7 with a KMD, where the Koopman mode associated with $g_{\varphi,k}$ is the dual product $\langle g_{\varphi,k}^{*}|g\rangle$. For example, if each observable in $\mathbf{g}=[g_{1},\ldots,g_{l}]^{T}$ lies in $\mathcal{S}$, we recover the vectorized form of Eq. 10 as an analogue of the classic KMD in Eq. 5:
$\mathbf{g}=\sum_{k\in\mathcal{I}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!\langle g_{\varphi,k}^{*}|\mathbf{g}\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi),\quad\text{so that}\quad\mathbf{g}=\sum_{k\in\mathcal{I}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!\exp(in\varphi)\langle g_{\varphi,k}^{*}|\mathbf{g}\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi).$
@@ 3 The Rigged DMD algorithm
(11) $g_{\theta}^{\epsilon}=\sum_{k\in\mathcal{I}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!K_{\epsilon}(\theta-\varphi)\,\langle g_{\varphi,k}^{*}|g\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi)\in L^{2}(\Omega,\omega),\qquad\text{where}\qquad g\in\mathcal{S}.$
This wave-packet approximates the part of Eq. 10 with spectral parameter $\lambda=\exp(i\theta)$ via convolution with $K_{\epsilon}$. The function $K_{\epsilon}$ is a highly localized smoothing kernel approximating a Dirac delta measure as $\epsilon\downarrow 0$. Therefore, the kernel filters out the contribution of unwanted modes to $g_{\theta}^{\epsilon}$ and amplifies the contribution of the desired modes at $\lambda=\exp(i\theta)$ for small $\epsilon$. In the limit as $\epsilon\downarrow 0$, we aim for $g_{\theta}^{\epsilon}$ to converge in a suitable sense (e.g., in $\mathcal{S}^{*}$) to a generalized eigenfunction of $\mathcal{K}$ associated with the point $\lambda=\exp(i\theta)$.
@@
(12) $K^{\rm(P)}_{\epsilon}(\varphi)=\frac{1}{2\pi}\frac{(1+\epsilon)^{2}-1}{1+(1+\epsilon)^{2}-2(1+\epsilon)\cos(\varphi)}$
be the Poisson kernel for the unit disc [50, p.16]. If we set $r=1+\epsilon$, then (see Section 5)
(13) $g_{\theta}^{\epsilon}=\frac{1}{4\pi}\left[(\mathcal{K}+r^{-1}e^{i\theta}I)(\mathcal{K}-r^{-1}e^{i\theta}I)^{-1}-(\mathcal{K}+re^{i\theta}I)(\mathcal{K}-re^{i\theta}I)^{-1}\right]g.$
@@ 3.1 Connections with previous work
Generalized eigenfunctions of Perron–Frobenius operators, which are pre-adjoints of Koopman operators, are well-studied. The decay rates of correlation functions, known as Pollicott–Ruelle resonances [68, 71], correspond to a distinct type of generalized eigenvalue of the Perron–Frobenius operator. Numerous studies have explored these expansions for specific examples of maps [4, 44, 72, 37, 45, 38, 79, 2, 3, 33, 1, 77]. A key distinction between these works and ours is that Pollicott–Ruelle resonances typically do not lie on the unit circle, corresponding instead to extensions of Koopman operators via meromorphic extensions of the resolvent. Operators that admit such extensions are termed Fredholm–Riesz operators [10]. In contrast, our work achieves a generalized eigenfunction expansion with eigenvalues on the unit circle and makes no assumption beyond $\mathcal{K}$ being unitary. Recently, Bandtlow, Just, and Slipantschuk [75, 11] and Wormell [88] have found a significant link in certain systems: EDMD eigenvalues in the large subspace limit can correspond to these resonances. Meanwhile, generalized eigenfunctions associated with the Koopman operator side have been less explored, with Mezić’s work on integrable Hamiltonian systems with one degree of freedom [64] being a notable exception. We revisit this example in Section 7.2, providing an explicit expansion for any number of degrees of freedom. A general-purpose algorithm for computing generalized eigenfunctions of Koopman operators from snapshot data in Eq. 2 has yet to be developed. Rigged DMD addresses this gap.
@@
The above papers compute spectral measures as opposed to generalized eigenfunctions. Nevertheless, we build on some of the above techniques. We use mpEDMD as a discretization method to construct data-driven approximations of the resolvent $(\mathcal{K}-zI)^{-1}$. Example 3.1 demonstrates the necessity for using mpEDMD. We then use the resolvent to compute smoothed approximations of generalized eigenfunctions. This smoothing is analogous to ResDMD, which computed smoothed approximations of spectral measures.
@@ 4.2 Measure-preserving EDMD
2:Compute an economy QR decomposition $\mathbf{W}^{1/2}\mathbf{\Psi}_{X}=\mathbf{Q}\mathbf{R}$, where $\mathbf{Q}\in\mathbb{C}^{M\times N},\mathbf{R}\in\mathbb{C}^{N\times N}$.

### v2

@@ Abstract
We introduce the Rigged Dynamic Mode Decomposition (Rigged DMD) algorithm, which computes generalized eigenfunction decompositions of Koopman operators. By considering the evolution of observables, Koopman operators transform complex nonlinear dynamics into a linear framework suitable for spectral analysis. While powerful, traditional Dynamic Mode Decomposition (DMD) techniques often struggle with continuous spectra. Rigged DMD addresses these challenges with a data-driven methodology that approximates the Koopman operator’s resolvent and its generalized eigenfunctions using snapshot data from the system’s evolution. At its core, Rigged DMD builds wave-packet approximations for generalized Koopman eigenfunctions and modes by integrating Measure-Preserving Extended Dynamic Mode Decomposition with high-order kernels for smoothing. This provides a robust decomposition encompassing both discrete and continuous spectral elements. We derive explicit high-order convergence theorems for generalized eigenfunctions and spectral measures. Additionally, we propose a novel framework for constructing rigged Hilbert spaces using time-delay embedding, significantly extending the algorithm’s applicability (Rigged DMD can be used with any rigging). We provide examples, including systems with a Lebesgue spectrum, integrable Hamiltonian systems, the Lorenz system, and a high-Reynolds number lid-driven flow in a two-dimensional square cavity, demonstrating Rigged DMD’s convergence, efficiency, and versatility. This work paves the way for future research and applications of decompositions with continuous spectra.
@@ 1 Introduction
This paper introduces a new Koopman mode decomposition and computational framework for unitary Koopman operators on a rigged Hilbert space (see Section 2.4). In the rigged Hilbert space, the discrete and continuous spectrum are treated on equal footing, and (almost) every point in the spectrum is associated with a set of coherent generalized eigenfunctions. [^1] The terminology is not to be confused with generalized eigenvectors that generate Jordan chains of a finite non-normal matrix [37]. Generalized eigenfunctions associated with the continuous spectrum are not elements of the Hilbert space. Instead, as distributions in the rigged Hilbert space, they represent coherent linear features of the observables that may be used to identify coherent patterns in Eq. 1.
Our algorithm, which we call Rigged DMD, computes smooth wave-packet approximations to generalized Koopman eigenfunctions using a two-stage framework (see Fig. 1). In the first stage, we use mpEDMD to construct a finite-dimensional unitary approximation to the Koopman operator from the snapshot data in Eq. 2. In the second stage, we use the mpEDMD resolvent to construct wave packets of observables with highly localized spectral content. These approximations converge to generalized eigenfunctions associated with both the discrete and the continuous spectrum (see Section 5). They are approximately coherent (see Theorem 5.13), achieve high orders of accuracy under natural regularity conditions on the spectral measure of $\mathcal{K}$ (see Theorem 5.9), and converge pointwise in regions where generalized eigenfunctions can be identified with smooth functions (see Theorem 5.11). For example, Figure 2 shows four wave-packet approximations to generalized eigenfunctions associated with the continuous spectrum of a Koopman operator on the Lorenz attractor. Rigged DMD provides the first general procedure for computing generalized eigenfunctions of Koopman operators.
There is extensive literature on rigged Hilbert spaces for self-adjoint operators in PDE analysis and quantum mechanics [49, 15, 58, 14, 13]. However, standard rigging techniques do not always apply to Koopman operators (consider the task of constructing a nuclear space on the Lorenz attractor). For this reason, we provide a novel method to construct nuclear spaces for unitary operators with sparse, infinite-dimensional matrix representations (see Section 6). This technique can be applied to any Koopman operator via time-delay embedding, enabling the computation of generalized Koopman eigenfunctions for the Lorenz attractor, for example. Moreover, Rigged DMD does not rely on time-delay embedding - it can be applied to any rigging, not necessarily constructed using time-delay embedding.
@@ 2 Background: Koopman mode decompositions
Such systems preserve an energy or volume associated with $\omega$ and are ubiquitous [7, 38, 84, 55, 98]. Many systems admit invariant measures [65] or measure-preserving post-transient behavior [71]. We also assume that the dynamical system is invertible modulo $\omega$-null sets (see [39, Chapter 7]), so that $\mathcal{K}$ is a unitary operator on the Hilbert space $L^{2}(\Omega,\omega)$:
@@ 2.2 Spectrum and pseudoeigenfunctions
Since $\mathcal{K}$ is unitary, we have $\mathrm{Sp}(\mathcal{K})\subset\mathbb{T}=\{z\in\mathbb{C}:|z|=1\}$. The spectrum includes the set of eigenvalues of $\mathcal{K}$, but in general, it may contain points that are not eigenvalues and hence do not have an associated eigenfunction. However, there are approximately coherent observables associated with each point because $\mathcal{K}$ is unitary (therefore, normal). For each point $\lambda\in\mathrm{Sp}(\mathcal{K})$ and any $\delta>0$, there exists a $\phi_{\delta}\in L^{2}(\Omega,\omega)$ with norm one such that $\|(\mathcal{K}-\lambda I)\phi_{\delta}\|\leq\delta$. The observable $\phi_{\delta}$ is called a $\delta$-pseudoeigenfunction [93, p. 31] and satisfies
@@ 2.3 Spectral measures and diagonalization
The spectral theorem [30, Thm. X.4.11] provides a diagonalization of $\mathcal{K}$ using a projection-valued measure $\mathcal{E}$ supported on $\mathrm{Sp}(\mathcal{K})\subset\mathbb{T}$. To each Borel subset $S\subset\mathbb{T}$, $\mathcal{E}$ associates an orthogonal projection $\mathcal{E}(S)$, which one should think of as the spectral content of $\mathcal{K}$ on $S$. In particular, we have the following decompositions:
@@ 2.4 Generalized eigenfunctions
To achieve a modal decomposition using generalized eigenfunctions, we require that $\mathcal{S}$ also be a countably Hilbert nuclear space [50, Ch. 1.3]. Then, there is a complete set of generalized eigenfunctions $\{g_{\theta,k}:\theta\in{[{-}\pi,\pi]_{\mathrm{per}}},k\in\mathcal{J}\}\subset\mathcal{S}^{*}$ such that for any $g\in\mathcal{S}$, it holds in $L^{2}(\Omega,\omega)$:
(10) $g=\sum_{k\in\mathcal{J}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!\langle g_{\varphi,k}^{*}|g\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi),\quad\mathcal{K}g=\sum_{k\in\mathcal{J}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!\exp(i\varphi)\langle g_{\varphi,k}^{*}|g\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi).$
The spectral measures $\{\xi_{g_{k}}\}$ reflect a choice of basis and normalization for the generalized eigenspaces. The multiplicity of $\lambda=\exp(i\theta)\in{\rm Sp}(\mathcal{K})$ is the number of nonzero functionals $\{g_{\theta,k}\}$ and the index set $\mathcal{J}$ is, in general, at most countable. This expansion replaces the projection-valued measure expansion in Eq. 7 with a KMD, where the Koopman mode associated with $g_{\varphi,k}$ is the dual product $\langle g_{\varphi,k}^{*}|g\rangle$. For example, if each observable in $\mathbf{g}=[g_{1},\ldots,g_{l}]^{T}$ lies in $\mathcal{S}$, we recover the vectorized form of Eq. 10 as an analogue of the classic KMD in Eq. 5:
$\mathbf{g}=\!\sum_{k\in\mathcal{J}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!\langle g_{\varphi,k}^{*}|\mathbf{g}\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi),\quad\!\text{so that}\!\quad\mathcal{K}^{n}\mathbf{g}=\!\sum_{k\in\mathcal{J}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!\exp(in\varphi)\langle g_{\varphi,k}^{*}|\mathbf{g}\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi).$
@@ 3 The Rigged DMD algorithm
(11) $g_{\theta}^{\epsilon}=\sum_{k\in\mathcal{J}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\!\!\!\!\!K_{\epsilon}(\theta-\varphi)\,\langle g_{\varphi,k}^{*}|g\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi)\in L^{2}(\Omega,\omega),\qquad\text{where}\qquad g\in\mathcal{S}.$
This wave-packet approximates the part of Eq. 10 with spectral parameter $\lambda=\exp(i\theta)$ via convolution with $K_{\epsilon}$. The function $K_{\epsilon}$ is a highly localized smoothing kernel approximating a Dirac delta measure as $\epsilon\downarrow 0$. Therefore, the kernel filters out the contribution of unwanted modes to $g_{\theta}^{\epsilon}$ and amplifies the contribution of the desired modes at $\lambda=\exp(i\theta)$ for small $\epsilon$. In the limit as $\epsilon\downarrow 0$, we aim for $g_{\theta}^{\epsilon}$ to converge in a suitable sense (e.g., in $\mathcal{S}^{*}$) to a generalized eigenfunction of $\mathcal{K}$ associated with the point $\lambda=\exp(i\theta)$. Note that while the generalized modes are distributions in the rigged Hilbert space and may not be associated with any standard functions in $L^{2}(\Omega,\omega)$ (for example, see Section 7.2.1), the wave-packet approximations in Eq. 11 are relatively well-behaved functions in $L^{2}(\Omega,\omega)$.
@@
(12) $K^{\rm(P)}_{\epsilon}(\varphi)=\frac{1}{2\pi}\frac{1-(1+\epsilon)^{-2}}{1+(1+\epsilon)^{-2}-2(1+\epsilon)^{-1}\cos(\varphi)}$
be the Poisson kernel for the unit disc [59, p.16]. If we set $r=1/(1+\epsilon)<1$, then (see Section 5)
(13) $g_{\theta}^{\epsilon}=\frac{1}{4\pi}\left[(\mathcal{K}+re^{i\theta}I)(\mathcal{K}-re^{i\theta}I)^{-1}-(\mathcal{K}+r^{-1}e^{i\theta}I)(\mathcal{K}-r^{-1}e^{i\theta}I)^{-1}\right]g.$
@@
  • There are three main parameters in the algorithm: $M$, $N$, and $\epsilon$. These parameters are often chosen adaptively in relation to one another. For a given dictionary and $N$, the error resulting from a finite $M$ corresponds to a quadrature error and can be controlled in various contexts [25, Section 4.1.3]. For a given $\epsilon$, we can determine a posteriori, through the computation of residuals and ResDMD [29, Appendix C], whether the dictionary is sufficiently rich to approximate the resolvent.
@@ 3.1 Connections with previous work
Generalized eigenfunctions of Perron–Frobenius operators, which are pre-adjoints of Koopman operators, are well-studied. The decay rates of correlation functions, known as Pollicott–Ruelle resonances [78, 81], correspond to a distinct type of generalized eigenvalue of the Perron–Frobenius operator. Numerous studies have explored these expansions for specific examples of maps [4, 53, 82, 46, 54, 47, 91, 2, 3, 36, 1, 89]. See also work on isolated spectra of quasi-compact transfer operators [60, 10, 35]. Froyland has developed numerical techniques for isolated spectra and eigenfunctions [41, 42, 43, 44, 32]. A key distinction between these works and ours is that Pollicott–Ruelle resonances typically do not lie on the unit circle, corresponding instead to extensions of Koopman operators via meromorphic extensions of the resolvent. Operators that admit such extensions are termed Fredholm–Riesz operators [11]. In contrast, our work achieves a generalized eigenfunction expansion with eigenvalues on the unit circle and makes no assumption beyond $\mathcal{K}$ being unitary. Recently, Bandtlow, Just, and Slipantschuk [87, 12] and Wormell [100] have found a significant link in certain systems: EDMD eigenvalues in the large subspace limit can correspond to these resonances. Meanwhile, generalized eigenfunctions associated with the Koopman operator side have been less explored, with Mezić’s work on integrable Hamiltonian systems with one degree of freedom [74] being a notable exception. We revisit this example in Section 7.2, providing an explicit expansion for any number of degrees of freedom using the Radon and Fourier transforms for the action and angle parts, respectively. A general-purpose algorithm for computing generalized eigenfunctions of Koopman operators from snapshot data in Eq. 2 has yet to be developed. Rigged DMD addresses this gap.
@@
The aforementioned papers compute spectral measures, but it is not always clear how their techniques can be extended to generalized eigenfunctions. Nevertheless, we build on some of these methods. We use mpEDMD as a discretization method to construct data-driven approximations of the resolvent $(\mathcal{K}-zI)^{-1}$. Example 3.1 demonstrates the necessity for using mpEDMD. We then use the resolvent to compute smoothed approximations of generalized eigenfunctions. This smoothing is analogous to ResDMD, which computed smoothed approximations of spectral measures.
@@ 4.2 Measure-preserving EDMD
2:Compute an economy pivoted QR decomposition $\mathbf{W}^{1/2}\mathbf{\Psi}_{X}=\mathbf{Q}\mathbf{R}$, where $\mathbf{Q}\in\mathbb{C}^{M\times N}$ and $\mathbf{R}\in\mathbb{C}^{N\times N}$.